### v2

# Is Your Imitation Learning Policy Better than Mine? Policy Comparison with Near-Optimal Stopping

Abstract
Imitation learning has enabled robots to perform complex, long-horizon tasks in challenging dexterous manipulation settings. As new methods are developed, they must be rigorously evaluated and compared against corresponding baselines through repeated evaluation trials. However, policy comparison is fundamentally constrained by a small feasible sample size (e.g., 10 or 50) due to significant human effort and limited inference throughput of policies. This paper proposes a novel statistical framework for rigorously comparing two policies in the small sample size regime. Prior work in statistical policy comparison relies on batch testing, which requires a fixed, pre-determined number of trials and lacks flexibility in adapting the sample size to the observed evaluation data. Furthermore, extending the test with additional trials risks inducing inadvertent p-hacking, undermining statistical assurances. In contrast, our proposed statistical test is sequential, allowing researchers to decide whether or not to run more trials based on intermediate results. This adaptively tailors the number of trials to the difficulty of the underlying comparison, saving significant time and effort without sacrificing probabilistic correctness. Extensive numerical simulation and real-world robot manipulation experiments show that our test achieves near-optimal stopping, letting researchers stop evaluation and make a decision in a near-minimal number of trials. Specifically, it reduces the number of evaluation trials by up to 37% as compared to state-of-the-art baselines, while preserving the probabilistic correctness and statistical power of the comparison. Moreover, our method is strongest in the most challenging comparison instances (requiring the most evaluation trials); in a multi-task comparison scenario, we save the evaluator more than 200 simulation rollouts.
Figure 1: Robot policy comparison problem under binary success/failure metrics. Novel policy $\pi_{1}$ is compared against baseline $\pi_{0}$ in a sequence of trials. Within a given evaluation budget, the evaluator seeks a statistically significant comparison in as few trials as possible. Due to the cost of hardware setup and calibration, as well as the limited inference speed of complex policies, these results generally arrive in sequence from a single (or few) hardware setups. Allowing the evaluator to adaptively and near-optimally tailor the number of trials based on the data observed so far — without compromising statistical assurances of the comparison — is a central contribution of this work.
I Introduction
The importance of trustworthy and efficient robot policy evaluation protocols has become paramount in imitation learning as the scale of underlying deep learning models and the complexity of tasks continue to increase. This need is especially pronounced in dexterous manipulation where stochastic, contact-rich interactions between the robot and the environment introduce inherent randomness in outcomes.
A particularly important aspect of policy evaluation is policy comparison, where two policies are repeatedly deployed in an environment to assess their relative performance. Policy comparison forms the foundation of robot learning as “an empirical science” [28], enabling researchers to objectively measure the scientific progress of the field. Nevertheless, this setting introduces an additional source of stochasticity due to random outcomes of both the first and the second policies, making the reliability of the comparison more challenging to ensure than evaluation of single policies.
To motivate concretely, consider an example policy comparison scenario presented in Fig. 1 where the performance is quantified based on binary success/failure metrics, a common choice [7, 59, 41] as continuous-valued rewards are often difficult to define. This scenario naturally arises when researchers want to demonstrate the effectiveness of a particular intervention (e.g., a new policy architecture) by comparing the new policy $\pi_{1}$ against a baseline $\pi_{0}$. Alternatively, $\pi_{1}$ and $\pi_{0}$ could represent the same policy evaluated under different environment distributions, providing insights on generalization. In either case, the evaluator faces two practical challenges. First, only a small number of trials (e.g., 10–60) [35, 18, 14, 40, 27, 5] can be performed per evaluation setting due to the large human effort needed to reset the environment between trials and the substantial wall-clock time imposed by limited inference throughput of large policies. While high-fidelity simulators can alleviate the human effort and still provide valuable insights into policy performance [32, 42], real-world evaluation remains indispensable for ensuring reliable deployment in downstream applications. Second, the evaluation results are revealed sequentially, possibly leading to fluctuating observations depending on when the evaluation is stopped. In the Fig. 1 example, the evaluator could observe more successes for $\pi_{0}$ after conducting additional trials, even though $\pi_{1}$ initially appeared superior after the first five.
Although recent work [53, 28] proposes statistical policy comparison approaches, it follows the conventional batch testing scheme, requiring complete results from a pre-determined number of trials before the statistical test can be performed. Furthermore, the test can be conducted only once for the corresponding results; even if the test fails to determine the relative performance due to closely matched results, the evaluator cannot append more evaluation trials to the existing results to run the test again, as doing so would constitute p-hacking that invalidates statistical assurances [43]. Unfortunately, this is a common but harmful research practice outside of robotics [23], which needs to be averted to ensure reproducible science.
To address these challenges, we propose a novel sequential testing framework named STEP (Sequential Testing for Efficient Policy Comparison) for rigorously comparing performance of imitation learning policies [^1] Although this paper focuses on imitation learning, STEP is naturally applicable to evaluating any types of policies based on binary metrics, including reinforcement learning (RL) policies with sparse 0/1 reward.. Unlike batch testing, our approach allows the number of trials to be varied within a given experimental budget. This critical feature offers two practical advantages. First, the evaluator can stop conducting trials early without sacrificing the probabilistic correctness of the comparison if enough statistical evidence is accumulated quickly in favor of $\pi_{1}$ (or $\pi_{0}$). Second, it reduces the epistemic risk of overconfident (and potentially incorrect) conclusions when $\pi_{0}$ and $\pi_{1}$ are closely matched, since the test will abstain from making a decision if statistical evidence remains low after all the planned trials have been executed. Alternatively, the evaluator may safely append additional trials to the original samples and re-conduct statistical analysis on all the results obtained thus far without inadvertent p-hacking. We demonstrate these advantages through simulation and real-world robot manipulation experiments. Furthermore, extensive numerical experiments show that STEP significantly outperforms state-of-the-art (SOTA) sequential methods, reducing the required number of trials by up to 37% without sacrificing probabilistic correctness. The specific contributions of this paper are as follows:
  • We propose STEP, a novel sequential statistical framework for evaluating relative performance of two policies with tunable probabilistic correctness  [^2] Associated website and code can be found at: https://tri-ml.github.io/step/.
  • Our sequential testing approach admits an adaptive number of evaluation trials tailored to the difficulty of the underlying comparison while achieving near-minimal sample complexity.
  • We additionally present a straightforward extension of our framework to (1) multi-task and (2) multi-policy comparison settings via a reduction to a set of pair-wise statistical comparisons.
II Preliminaries
Consider a physical robot trained to complete a task in a variety of environment realizations. This setting is naturally modeled as a partially observable Markov decision process (POMDP) [24], where the underlying state $s$ represents the environment and robot states. The observation $o$ is determined by the robot’s embodiment and sensing apparatus. The training pipeline is designed to synthesize a policy taking actions $a$ that achieve a high reward $r(s,a)=\mathbbm{1}[s\in\mathcal{S}_{\text{success}}]$ on a particular task, where $\mathcal{S}_{\text{success}}$ is a set of successful terminal states in which the episode will terminate. This reward encodes a binary success/failure criterion of the task. In imitation learning, a surrogate loss function is used to train a policy that matches the behavior of expert demonstrations [46] instead of directly solving the POMDP. At the end of a training process, a policy $\pi_{1}$ will be obtained. This policy has a true success rate (i.e., expected total episode reward) under a distribution $\mathcal{D}_{s_{0},o_{0}}$ over the initial state and observation:
$p_{1}=\mathbb{E}_{\mathcal{D}_{s_{0},o_{0}},\pi_{1}}\bigg{[}\sum_{t=0}^{T}r(s_{t},a_{t})\bigg{]},$ (1)
where dependence on the state transition and observation models are omitted for brevity. The true success rate is unknown and must be estimated via multiple evaluation trials.
Assumption 1 (Regularity)
In each evaluation trial, the initial state $s_{0}$ and the observation $o_{0}$ are drawn in an independent and identically distributed (i.i.d.) fashion from the underlying distribution $\mathcal{D}_{s_{0},o_{0}}$ [^3] Note that this is a standard assumption in statistical testing. A discussion of practical methods by which to approximately satisfy this condition in robotic evaluation is included in [28].. We assume access to samples from $\mathcal{D}_{s_{0},o_{0}}$, but do not assume $\mathcal{D}_{s_{0},o_{0}}$ itself to be known.
Under Assumption 1, the $n^{\text{th}}$ evaluation trial involves making an i.i.d. draw of an environment from $\mathcal{D}_{s_{0},o_{0}}$ and running the policy $\pi_{1}$ in this environment. This yields a binary success/failure outcome $z_{1,n}$ corresponding to an i.i.d. draw from a Bernoulli random variable with mean $p_{1}$, which is the true performance (success rate) of the policy $\pi_{1}$ on the task:
$z_{1,n}\sim\text{Ber}(p_{1}).$ (2)
Here, $z_{1,n}=1$ indicates success and $z_{1,n}=0$ failure. For any baseline policy $\pi_{0}$, we similarly denote the outcome of its $n^{\text{th}}$ trial as $z_{0,n}\sim\text{Ber}(p_{0})$. For the sake of statistical analysis, we pair the outcomes of two policies by their indices in a vector $Z_{n}=(z_{0,n},z_{1,n})$. The policy comparison problem can be formalized in the sense of Neyman-Pearson (N-P) statistical testing [37], where the null (skeptical) hypothesis is that the novel policy $\pi_{1}$ is no better than the baseline $\pi_{0}$ and the alternative is that the novel policy is indeed better:
Null Hypothesis $\begin{split}\text{Null Hypothesis }\mathbb{H}_{0}:&\hskip 5.69054ptp_{1}~{}\leq~{}p_{0}\equiv(p_{0},p_{1})\in\mathcal{H}_{0}\\ \text{Alt. Hypothesis }\mathbb{H}_{1}:&\hskip 5.69054ptp_{1}~{}>~{}p_{0}\equiv(p_{0},p_{1})\in\mathcal{H}_{1}.\end{split}$ (3)
Figure 2: The policy comparison problem as a composite-vs-composite statistical test. The null hypothesis set (red) corresponds to the novel policy being worse ($p_{1}~{}\leq~{}p_{0}$), while the alternative hypothesis set (blue) corresponds to the novel policy being better ($p_{1}>p_{0}$). The sets are each termed “composite” because they contain many elements. For any pair of policies, the truth corresponds to a single point; as examples, the white circle is a case where the alternative is true (baseline success 45%, novel policy success 55%), while the black star corresponds to the null being true (baseline success 82%, novel policy success 35%).
As illustrated in Fig. 2, this amounts to a composite vs. composite statistical test [4]. A hypothesis is termed “simple” if it is singleton, i.e., corresponds to a single data-generating distribution. For example, $(p_{0},p_{1})=(0.3,0.7)$ would be a simple hypothesis. If the hypothesis corresponds to multiple data-generating distributions, it is termed “composite”.
The Type-I error rate (denoted $\alpha$) of a statistical test corresponds to the probability of falsely rejecting the null under the worst-case singleton element $h_{0}$ of the null hypothesis set $\mathcal{H}_{0}$ – that is, under the hardest-to-distinguish $(p_{0},p_{1})\in\mathcal{H}_{0}$. This type of error corresponds to falsely concluding that $\pi_{1}$ is better than $\pi_{0}$ when it is not. The power of a test (denoted $1-\beta$) is the mixture probability of correctly rejecting the null when the alternative is true, under some measure on the set of alternatives [^4] For example, a minimax measure mimics the Type-I setting (e.g., the worst-case singleton $h_{1}\in\mathcal{H}_{1}$). However, due to the finite termination condition, such a measure is vacuous because there will always exist a $(p_{0},p_{0}+\epsilon)$ for $\epsilon>0$ sufficiently small such that the attainable power will only negligibly exceed random guessing. Therefore, in practice other measures must be used.. The Type-II error $\beta$ is the associated (mixture) probability of failing to reject the null when the alternative is true. This represents failing to conclude that $\pi_{1}$ is better, when it is in fact better than $\pi_{0}$.
Finally, while the preceding development has focused on one-sided testing, the comparison problem naturally admits a bidirectional version allowing decisions for membership in either the alternative set (RejectNull) or the null set (AcceptNull). Note that the decision AcceptNull formally amounts to rejecting the null of a “flipped” version of Eq. 3. In all subsequent discussion we will implicitly utilize the bidirectional test, which will allow for the decision AcceptNull.
III Problem Formulation
We assume that a robot evaluator is tasked with distinguishing two policies via successive evaluations, resulting in the testing paradigm described in Section II. We also assume that the evaluator has pre-selected the desired significance level $\alpha^{*}$ of the comparison and a maximum number of trials (for each policy) that they are willing or able to run: $N_{\text{max}}$.
First, we note that the Type-I error must always be controlled in statistical testing, i.e., upper-bounded at the evaluator-specified rate $\alpha^{*}\in(0,1)$. This represents a hard constraint (“validity”) that will be enforced in all subsequent testing procedures; a test is not feasible if it is not Type-I error controlling. The evaluator’s goal is to synthesize a decision rule that limits the Type-I Error to $\alpha^{*}$ while maximizing power and minimizing the expected number of evaluation trials.
We assume an underlying representation of the evaluation information collected thus far is available to the evaluator in a state $x_{n}=F(Z_{1},Z_{2},...,Z_{n})$. Then, the evaluator’s procedure consists of deciding whether to Continue (gather another trial for each policy) or stop (and either AcceptNull or RejectNull). Given a decision set $\mathcal{U}=\{\textbf{AcceptNull},\textbf{Continue},\textbf{RejectNull}\}$, the problem is to find a state partition $\zeta=u(x)$ to optimally balance minimizing the expected sample size and maintaining high power, conditioned on the Type-I Error rate constraint, as shown in Eq. 4:
$\begin{split}\min_{\zeta:\mathcal{X}\mapsto\mathcal{U}}&\hskip 5.69054pt\mathbb{E}_{\mu(\mathcal{H}_{1})}[n_{\text{stop}}+c\beta_{N_{\max}}]\\ \text{s.t. }&\hskip 5.69054pt\max_{h_{0}=(p_{0},\text{ }p_{1})\in\mathcal{H}_{0}}\alpha(\zeta,h_{0})~{}\leq~{}\alpha^{*}\\ &\hskip 5.69054pt0~{}\leq~{}n_{\text{stop}}~{}\leq~{}N_{\text{max}}\text{ w.p.1}.\end{split}$ (4)
This function is a multi-objective optimization which seeks to simultaneously minimize expected sample size and maximize power subject to the validity constraint. Informally, for any feasible terminal power $1-\beta_{N_{\max}}^{\text{feasible}}$ (conditioned on $N_{\text{max}},\alpha^{*}$, and the true underlying distribution, which we emphasize is unknown a priori), there is a value of $c>0$ that effects a decision rule in Eq. 4 approximating a test controlling $\beta=\beta_{\text{feasible}}$. For example, this framework recovers the batch problem as $c\rightarrow\infty$ (demanding maximal power), and immediate termination as $c\rightarrow 0^{+}$ [^5] For the latter case: without looking at any data, draw a random number uniformly on [0, 1]. If it is less than $\alpha^{*}$ reject the null, otherwise fail to reject and terminate. This terminates at step 0 with probability 1, has power $\alpha^{*}$, and is valid. (demanding minimal sample complexity). This objective will govern the methodology and analysis presented in the rest of the paper.
IV Methodology
There are several important practical considerations to constructing a near-optimal solution to Eq. 4. We present the concrete challenges first, and then discuss the technical innovations that account for them. Throughout, we will leverage significant mathematical structure in the testing problem. Where insightful or intuitive, this will be explained in situ. See Section X-F in the Supplement for additional details.
IV-A State Representation
First, we express the information available to the evaluator at time $n$ (i.e., after $n$ evaluation trials have been performed for each policy) in a control-theoretic state representation; the evaluator’s decision can then be understood as a state-feedback decision rule. In selecting this representation, the first instance of mathematical structure is the membership of Bernoulli distributions in the univariate exponential family; such distributions have known sufficient statistics which represent (informally) an optimal compression of the data for the purposes of testing and estimation [4]. Thus, a natural (“near-minimal”) state representation is precisely the element-wise sufficient statistic for $p_{0}$ and $p_{1}$ respectively, augmented with a time state. For univariate exponential family distributions, the sufficient statistic is the sum of the observed data (i.e., respective number of successful trials under $\pi_{0}$ and $\pi_{1}$); this makes the state representation a first-order discrete integrator, as shown in Eq. 5:
$\begin{split}\mathbf{x}_{n+1}&=\mathbf{x}_{n}+\mathbf{d}_{n}\\ \mathbf{d}_{n}&=(z_{0,n},z_{1,n},1)\sim(\text{Ber}(p_{0}),\text{ }\text{Ber}(p_{1}),\text{ }1)\\ \mathbf{x}_{0}&=\mathbf{0}_{3}.\end{split}$ (5)
Unlike in a typical control problem, we cannot actively guide the state trajectory; we are instead deciding when to stop based on the state trajectory. Concretely, the “control” involves partitioning the state space into stopping and continuation regions. In robotics, a similar set-theoretic notion arises in robust control and safe navigation, through the generation of invariant sets for dynamic systems (i.e., in reachability and barrier function-type methods [48, 2]), though those methods are primarily interested in nonstochastic uncertainty. Separately, many applications of asset pricing in mathematical finance utilize optimal stopping to set options prices under stochastic uncertainty [36].
IV-B Decision Regions
In the sequential problem the state space partitions occur in sequence, yielding an effective representation of the decision rule $\zeta:\mathcal{X}\mapsto\mathcal{U}$ in the form of Eq. 6:
$\zeta\equiv\bigg{\{}\mathcal{X}^{\text{Reject Null}}_{n},\mathcal{X}^{\text{Accept Null}}_{n},\mathcal{X}^{\text{Continue}}_{n}\bigg{\}}_{n=1}^{N_{\text{max}}}.$ (6)
At each time step, the sets jointly encode control decisions for every state. Intuitively, the larger the size of the rejection region $\mathcal{X}_{n}^{\text{Reject Null}}$ (resp. $\mathcal{X}_{n}^{\text{Accept Null}}$), the smaller the number of expected trials needed to reject the null (resp. alternative), achieving a lower value of the stopping time component of Eq. 4. Focusing on $\mathcal{X}_{n}^{\text{Reject Null}}$ in the following development [^6] As discussed in Section II, the case of rejecting the alternative and accepting the null can be considered by flipping $p_{0}$ and $p_{1}$., our auxiliary objective is then to maximize the size of the set $\mathcal{X}_{n}^{\text{Reject Null}}$ globally across all $n\in\{1,\cdots,N_{\max}\}$. We take a probabilistic approach, allowing the states to reject with a probability less than 1. As we will see in Section IV-E, this yields an efficient optimization problem.
In addition to the maximization, there are two core challenges to synthesizing such a partition. First, the Type-I Error rate must be controlled. The decision to stop and reject the null hypothesis in a state $\mathbf{x}_{n}$ incurs risk due to the probability that $\mathbf{x}_{n}$ was reached under data generated from some $(p_{0},p_{1})\in\mathcal{H}_{0}$ – i.e., under the null hypothesis in Eq. 3. Thus, having a large $\mathcal{X}^{\text{Reject Null}}_{n}$ risks violating Type-I Error rate control. Second, the temporal rate at which the risk is accrued must be set appropriately to encourage early stopping on easy instances (i.e., $\pi_{1}$ significantly outperforms $\pi_{0}$) without harming performance too much on harder instances (i.e., $\pi_{1}$ and $\pi_{0}$ are closely competing, which requires many trials to distinguish). We will first address Type-I Error control in Section IV-C, and the temporal risk accumulation in Section IV-D. The resulting tractable optimization problem is presented in Section IV-E.
IV-C Type-I Error Control
The composite nature of the null hypothesis (which contains all pairs $\mathcal{H}_{0}=\{(p_{0},p_{1})\in[0,1]^{2}\text{ }\rvert\text{ }p_{0}~{}\geq~{}p_{1}\}$) means that the decision-making problem can be thought of (informally) as analogous to distributionally robust control, where the uncertainty is over the particular worst-case $(p_{0},p_{1})\in\mathcal{H}_{0}$. Controlling the Type-I Error in policy comparison then amounts to controlling the Type-I Error uniformly (robustly) for all $h\in\mathcal{H}_{0}$. Suppose that some rejection region for the first $n-1$ steps has been obtained with accumulated risk $\alpha_{n-1}$, and we are interested in bounding the Type-I Error for $n$ by some $\alpha_{n}>\alpha_{n-1}$. Given the notion of stopping and continuation regions introduced in the previous section, this is expressed mathematically as:
$\max_{h\in\mathcal{H}_{0}}\mathbb{P}_{h}\left(\mathbf{x}_{n}\in\mathcal{X}^{\text{Reject Null}}_{n}\mid\mathcal{X}_{n-1}^{\text{Reject Null}}\right)~{}\leq~{}\alpha_{n}.$ (7)
The dependence on the (probabilistic) rejection region for $n-1$ is made explicit in Eq. 7, reflecting the internal dynamic structure. For example, if some state $\mathbf{x}_{n-1}=(a,b,n-1)$ is rejected at $n-1$ with a non-zero probability, then the 1-step reachable states (e.g., $(a+1,b,n)$) under the dynamics Eq. 5 are less likely to be feasible at $n$. The presence of many (infinite) elements $h$ in the set of null hypotheses $\mathcal{H}_{0}$ makes verification of Type-I Error control challenging a priori, even for a single particular $n$. However: monotonicity, symmetry, and smoothness properties of this family of testing problems (noted in, for example, [13]) allow for efficient discretization procedures that preserve safety. Specifically, it suffices to consider a discrete set of the “worst-case” nulls ^$\hat{\mathcal{H}}_{0}=\{(p^{(1)},p^{(1)}),(p^{(2)},p^{(2)}),\cdots,(p^{(M)},p^{(M)})\}$, where $10~{}\leq~{}p^{(1)}\cdots~{}\leq~{}p^{(M)}~{}\leq~{}$ 1. Further details of this discretization process is given in Section X-F2 in the Supplement. Thus, Eq. 7 reduces to a finite set of individual inequality constraints.
We can equivalently represent this set of constraints as a linear inequality $P_{n}\mathbf{w}_{n}~{}\leq~{}\alpha_{n}\mathbbm{1}$, where $\mathbf{w}_{n}$ is a vector representing the probability of rejecting the null in each state $\mathbf{x}_{n}\in\mathcal{X}_{n}=\{(0,0,n),(0,1,n),\cdots(N_{\max},N_{\max},n)\}$. $P_{n}$ is a non-negative matrix of size $(M,|\mathcal{X}_{n}|)$ where each row represents the probability of reaching particular states under $h^{(i)}=(p^{(i)},p^{(i)})$:
$(P_{n})_{ij}=\mathbb{P}_{h^{(i)}}\left(\mathbf{x}_{n}=\mathbf{x}_{n}^{j}\mid\mathcal{X}_{n-1}^{\text{Reject Null}}\right),$ (8)
where $\mathbf{x}_{n}^{j}$ is the $j^{\text{th}}$ (discrete) state in $\mathcal{X}_{n}$. Given the rejection region from the previous time step $n-1$, we can accurately compute this probability Eq. 8 by forward-propagating the previous state occupancy distribution $(P_{n-1})_{i}$ according to the stochastic dynamics model Eq. 5 under $h^{(i)}$.
IV-D Power Adaptivity to Varying Difficulty: Risk Budgets
As discussed in Section IV-B, we must appropriately adjust the temporal rate of risk accumulation. To formally define this notion of risk, we introduce a non-negative scalar function $f(n)$ for $n\in\{1,\cdots,N_{\max}\}$, which determines the maximum allowable Type-I Error under any null hypothesis at each step $n$. We impose a constraint $\sum_{n=0}^{N_{\text{max}}}f(n)=\alpha^{*}$ to globally bound the Type-I error by $\alpha^{*}$. This risk “budget” can be interpreted as encoding the evaluator’s competing objectives: to reject the null and stop the evaluation quickly in easier cases (front-loading the risk accumulation) against the desire to wait longer to achieve a significant decision in harder instances (delaying risk accumulation until more data is collected). Importantly, any risk budget $f$ that is nonnegative everywhere and sums to some $r\in[0,\alpha^{*}]$ maintains Type-I error control at level $\alpha^{*}$. However, the shape of the risk budget will significantly influence the power of the resulting procedure, and thus represents an important component in solving (near-optimally) Eq. 4. With this said, in the following experiments we fix the risk budget to be uniform in order to focus on optimizing the decision regions (described in Section IV-E); to be explicit: this means the risk budget is $f(n)=\frac{\alpha^{*}}{N_{\text{max}}}$ for each scenario. This selection is heuristically reasonable, but leaves potential for further improvements in future work.
IV-E Tractable Optimization
Having specified the risk budget $f(n)$, it is straightforward to verify that the Type-I Error control is achieved if the following constraint is satisfied for all $n\in\{1,\cdots,N_{\max}\}$:
$P_{n}\mathbf{w}_{n}~{}\leq~{}\sum_{k=1}^{n}f(k)\mathbbm{1}$ (9)
Under this constraint, our objective is to maximize the size of the rejection region. We propose to solve the following series of optimization problems to tractably construct the rejection regions, one for each $n$:
[…]
This objective encourages the rejection from as many states as possible, maximizing the size of $\mathcal{X}_{n}^{\text{Reject Null}}$. Furthermore, it implicitly rejects from states less likely to occur under any null hypothesis, which are “cheaper” in terms of accruing risk. The first constraint ensures that the Type-I error is controlled up to time $n$, as discussed in Section IV-C and Section IV-D. The second constraint is to enforce boundedness of rejection probabilities in $[0,1]$.
Algorithm 1 STEP Decision-Rule Synthesis
  Input: $N_{\text{max}}>0$, risk budget function $f(n)$, type-I error limit $\alpha^{*}\in(0,1)$, number of approximation points $M$
  Initialize: $\zeta_{0}=\emptyset$, $(P_{0})_{ij}=1$ if $(i,j)=(0,0)$ else 00.
  for n $\in\{1,...,N_{\text{max}}\}$ do
     $P_{n}\leftarrow\textbf{Propagate}(P_{n-1},\zeta_{n-1},M)$Eq. 8}
     $\mathbf{w}_{n}\leftarrow\textbf{Opt}(P_{n},f)$Eq. 10}
     $\zeta_{n}\leftarrow\textbf{Compress}(\mathbf{w}_{n})$
  end for
  return  $\zeta=\{\zeta_{1},\dots,\zeta_{N_{\text{max}}}\}$STEP policy}
The optimization problem Eq. 10 is a simple linear program that can be efficiently solved by a standard optimization software; it is a maximization of $\lVert\mathbf{w}_{n}\rVert_{1}$ over the nonnegative orthant, subject to additional linear inequality constraints to control Type-I Error. Because each $P_{n}$ depends on the corresponding $\zeta_{n-1}$, the optimization needs to be performed sequentially for each $n$ in increasing order. Nevertheless, all the computation can be performed offline prior to running the actual policy evaluation. See Algorithm 1 for the decision-rule synthesis procedure. The fact that the optimization problems are sequentially solvable is owed to the isolation of the risk accumulation rate $f(n)$ as a tunable parameter; otherwise, the rejection regions would be generally coupled across $n$ and the optimization would be more challenging. Finally, we note in Algorithm 1 that the representation of the optimal policy can be compressed substantially from the vector $\mathbf{w}_{n}$ to a set of connected sets $\zeta_{n}=\bigg{\{}\mathcal{X}^{\text{Reject Null}}_{n},\mathcal{X}^{\text{Accept Null}}_{n},\mathcal{X}^{\text{Continue}}_{n}\bigg{\}}$ defined by their boundaries, due primarily to the general monotonicity of the optimal regions and the unimodality of the distribution of the test statistic.
V Experiments
Figure 3: False positive rate of four feasible methods (Barnard, SAVI, Lai, and Ours (STEP)) and the SPRT (Oracle method) for 400 simulated trajectories drawn from the respective worst-case null distributions for each of 45 alternatives (squares in color); $N_{\text{max}}=500$ and $\alpha^{*}=0.05$. Note that naively utilizing a batch method in sequence leads to violation of Type-1 Error control (red). Additionally, note that SAVI and Lai struggle to utilize the full risk budget in finite $N_{\text{max}}$ (darker blue regions).
The experiments are designed to investigate the following key aspects of effective evaluation:
  1. A.
    How does sequential testing compare with batch testing in terms of statistical validity (Type-1 Error)?
  2. B.
    What are the unique advantages of STEP in sequential comparison problems of varying difficulty?
  3. C.
    How sample efficient is STEP in practical policy comparison settings?
The experiments address these questions through extensive numerical validation on simulated success/failure data, as well as practical validation on both simulated rollouts and tasks on physical hardware.
V-A Baseline Procedures
The baselines constitute the SOTA sequential analysis methods described in Lai [30] and Lai and Zhang [31] (termed “Lai”) and the Safe, Anytime-Valid Inference (SAVI) method of Turner and Grünwald [50], which is specifically tailored to contingency table (i.e., policy comparison) problems (termed “SAVI”). Similar to STEP, Lai is a valid sequential method under a pre-determined $N_{\max}$ and is asymptotically optimal as $N_{\max}$ tends to infinity. The specific difference between these methods is that Lai makes the asymptotic assumption of normality of the empirical means. This allows for any distribution subject to the central limit theorem (CLT) to be analyzed tractably in the large-data regime, as the resulting optimal stopping PDE amounts to the solution of a reverse heat equation. STEP, by contrast, solves the exact PDE at the fundamental finite-sample resolution of the testing problem. As will be shown, the Lai procedure does best in difficult (i.e., requiring a lot of data) and symmetric (i.e., $p_{0},p_{1}$ each near $0.5$) test settings, as these best match the asymptotic approximation where the CLT approximation is well-matched to the exact distribution of the empirical means. SAVI, by contrast, does not require a fixed $N_{\max}$ a priori and maintains validity for an arbitrarily large $N_{\max}$. In essence, SAVI procedures construct a test statistic subject to stochastic dynamics (the sequential incorporation of data) which possesses a high-probability ‘stability certificate’ under any null data generating process. The ‘stability’ in this case grants time-uniformity (in perpetuity, i.e., not requiring any finite $N_{\text{max}}$); thus, under any null hypothesis, the probability of ever violating the certificate is bounded (by design) to be less than $\alpha$. In exchange for increased generality and robustness, however, SAVI methods inevitably sacrifice sample complexity in cases where a finite upper-bound to $N_{\max}$ is determined by practical time and resource constraints that researchers are subject to. We emphasize that each existing method is widely applicable to practical testing problems beyond robotics, as is STEP. Importantly, STEP is tailored specifically to the small-$N_{\text{max}}$ evaluation regime.
Additionally, an Oracle Sequential Probability Ratio Test (SPRT) [55] is run using the true singleton alternative and associated worst-case singleton null; this method represents a near-optimal procedure for easier problems where feasible methods quickly approach a terminal power close to 1 and still serves as a reasonable benchmark in harder problems. We emphasize that in practical cases, the Oracle method is infeasible to the evaluator; it is included to give a conservative estimate of the optimality gap of each method. Additional information about each baseline is included in Section VI.
Figure 4: Terminal power of three feasible methods (SAVI, Lai, and Ours (STEP)) and the SPRT (Oracle method) for 5000 simulated trajectories on each of 45 alternatives (squares in color); $N_{\text{max}}=500$ and $\alpha^{*}=0.05$. Because $N_{\text{max}}$ is large, the terminal power is high for all but the most difficult cases; however, note that SAVI has significantly worse power in these difficult instances where $p_{0}$ and $p_{1}$ are closely competing. This is due to its inherent validity for arbitrary $N_{\text{max}}$, which is unnecessary in modern robotics evaluation contexts and renders the methods conservative in harder instances.
V-B Numerical Simulation
For evaluation, we discretize a grid over the space of alternatives in 10-percentage point increments, offset from zero by five points. There are forty-five resulting alternatives in which $p_{1}>p_{0}$: $(p_{0},p_{1})=\{(0.05,0.15),(0.05,0.25),...,(0.85,0.95)\}$. For each of these alternatives, 5000 trajectories (each of length $N_{\text{max}}$) are simulated from $(\text{Ber}(p_{0}),\text{Ber}(p_{1}))$. In addition, 400 additional trajectories are generated under the worst-case null for each of the 45 alternatives, in order to verify the Type-I Error control. This evaluation data is shared across each methodology.
The algorithms are first validated on multiple pairs of $(N_{\text{max}},\alpha^{*})$. The ensuing numerical results will use the case $N_{\text{max}}=500,\alpha^{*}=0.05$; similar figures for $N_{\text{max}}=100,\alpha^{*}=0.05$ are included in Section X-B in the Supplement. For each method, we compute for each of the 45 alternatives the following characteristics: (1) Type-I Error – the fraction of associated worst-case null trajectories which have incorrectly rejected the null; (2) Terminal Power – the fraction of alternative trajectories which have correctly rejected the null by step $N_{\text{max}}$; (3) Cumulative Power – a visualization of the fraction of alternative trajectories which have correctly rejected the null by step $t$ for all $t\in\{0,1,...,N_{\text{max}}\}$. Note that another natural evaluation metric, the expected time-to-decision ($\mathbb{E}[n_{\text{stop}}]$), can be derived from the cumulative power as the area between the curve and the constant $y=1$.
V-B1 How does sequential testing compare with batch testing in Type-I Error control?
We show the Type-1 Error control of each sequential method in Fig. 3, as well as a widely-used batch method (Barnard’s Test [3]) run in sequence. Each sequential method maintains Type-1 Error control; however, STEP is the most efficient at using the full risk budget (lightest blue), while Lai is weaker (more conservative) in the lower-variance regime and SAVI is weaker in general (darker blue). Using a batch method like Barnard’s Test in sequence, conversely, violates Type-1 Error control (red). This highlights the aforementioned p-hacking issue of the batch testing scheme. A rigorous batch evaluator, having chosen $N$, cannot adapt to the data as it arrives lest they invalidate a resulting conclusion. On the other hand, STEP provides a safety margin for continued testing with the maximum allowable sample size $N_{\max}>N$.
V-B2 What are the unique advantages of STEP in comparison problems of varying difficulty?
Terminal Power
We begin a more fine-grained comparison of the efficiency of sequential procedures by presenting the terminal power (probability of deciding RejectNull by step $N_{\text{max}}$) of all feasible methods and the Oracle SPRT in Fig. 4. This metric serves to illustrate a downside of SAVI: its inherent validity at arbitrarily large $N_{\text{max}}$ imposes strong finite-time costs. In this case, the terminal power when the true gap in policy performance is 10 percentage points is significantly lower than the Lai baseline and our STEP, which closely approximate the SPRT Oracle.
Cumulative Power
We now demonstrate the downside of the Lai baseline procedure as compared to our method through the cumulative power of the procedure.
In Fig. 5 (right), we illustrate the cumulative power on a hard evaluation case. This represents a second view of the observation presented in the terminal power analysis: the Lai procedure and our STEP each significantly outperform the SAVI procedure in the small-gap regime.
Conversely, in Fig. 5 (center), we illustrate a difficult setting in which the Lai procedure struggles. In this setting the performance gap is again 10 percentage points, but the distribution is lower-variance and skewed as compared to that of Fig. 5 (right); as such, the Lai procedure cannot effectively adapt and our STEP significantly outperforms in terms of deciding more quickly for the alternative.
Figure 5: Cumulative power of all feasible methods (Lai, SAVI, STEP (Ours)) and SPRT Oracle over 5000 trajectories in three evaluation settings of increasing difficulty; $(p_{0},p_{1})$ for each setting title the respective figures. $N_{\text{max}}=500$ and $\alpha^{*}=0.05$. The expected time-to-decision is the integral of the area above the cumulative power curve; therefore, curves higher and to the left are better (black arrow). (Left) For a gap of 30 percentage points, all methods demonstrate similar stopping times. (Center) For a gap of 10 percentage points in the low-variance regime (i.e., farther from 0.5), STEP significantly outperforms the Lai procedure and is better than SAVI. (Right) For a gap of 10 percentage points in the high-variance regime, SAVI struggles to maintain high power, and underperforms the other methods.
V-C How sample efficient is STEP in practical policy comparison settings?
In addition to the numerical validation, we evaluate STEP through two sets of real-world robot evaluation experiments and a simulation experiment. In the first set of experiments, we compared policies with noticeable performance gaps to demonstrate the early-stopping capability of our approach. In the second set, we compared closely-competing policies to characterize necessary sample sizes for statistical validity when policy performance gaps become small. In the simulation experiment, we perform multi-task comparison of two SOTA imitation learning policies. See Section X-A in the Supplement for more details on the hardware experiments.
V-C1 Hardware Evaluation in the Large/Medium-Gap Regime
In this experiment, we consider two manipulation tasks for a bimanual Franka Emika Panda robot: FoldRedTowel (Fig. 9) and CleanUpSpill (Fig. 9 and Fig. 9). We trained single-task diffusion policies [14] on each task, with 300 human demonstrations for FoldRedTowel and 150 for CleanUpSpill, respectively. In addition to the RGB images, the policy receives the proprioceptive states as additional observations.
In FoldRedTowel, we compare two policy checkpoints from a single training run. The baseline policy $\pi_{0}$ was trained for 10000 gradient steps with an AdamW [34] optimizer and achieved the action mean-squared error (MSE) of 1.61E-3 on the validation set. The other policy $\pi_{1}$ continued the training for 70000 additional steps, yielding the validation action MSE of 1.35E-3. To evaluate each policy, five in-distribution (ID) initial conditions were chosen and repeated 10 times each, constituting 50 total trials. ID initial conditions visually resemble the ones that appear in the training dataset. As shown in Table I (rows 1-3), the empirical gap in success rates was 36 percentage points (56% to 92% success), suggesting that $\pi_{0}$ was under-trained. Each sequential method detected a significant difference at level $\alpha^{*}=0.05$ in 19 – 23 trials (apiece). That is, had the algorithms been active during collection, the last 27 – 31 rollouts per policy were unnecessary for confirming the improvement of the later checkpoint over the former. Additionally, note that the Lai method and our STEP sequential procedures were each tuned for an $N_{\text{max}}$ of 50 (row 1), 200 (row 2), and 500 (row 3) rollouts. In the latter two cases, additional rollouts could have been run up to 200 or 500 per policy if the gap was smaller without compromising the validity of the decision.
Task Type $\alpha^{*}$ $N_{\text{max}}$ $N$ $\hat{p}_{0}$ $\hat{p}_{1}$ SAVI Lai STEP (Ours) SPRT∗∗∗
FoldRedTowel $\pi_{i}$ 0.05 50 50 0.560 0.920 20 19 20 17
FoldRedTowel $\pi_{i}$ 0.05 200 50 0.560 0.920 20 20 21 17
FoldRedTowel $\pi_{i}$ 0.05 500 50 0.560 0.920 20 23 23 17
CleanUpSpill $\mathcal{D}^{i}_{s_{0},o_{0}}$ 0.05 50 50 0.280 0.800 7 9 10 7
CleanUpSpill $\mathcal{D}^{i}_{s_{0},o_{0}}$ 0.05 200 50 0.280 0.800 7 12 10 7
CleanUpSpill $\mathcal{D}^{i}_{s_{0},o_{0}}$ 0.05 500 50 0.280 0.800 7 14 14 7
CarrotOnPlate $\mathcal{D}^{i}_{s_{0},o_{0}}$ 0.05 100 100 0.590 0.680
CarrotOnPlate $\pi_{i}$ 0.05 100 100 0.680 0.760
SpoonOnTowel $\pi_{i}$ 0.01 500 500 0.084 0.386 33 35 33 26
EggplantInBasket $\pi_{i}$ 0.01 500 500 0.400 0.564 192 123 119 128
StackCube $\pi_{i}$ 0.01 500 500 0.000 0.030 329 403 199 135
Multitask $\pi_{i}$ 0.03 1500 1500 N/A N/A 554 561 351 289
TABLE I: Empirical time-to-correct-decision for all hardware (top) and simulation (bottom) policy comparisons. The comparison type is described first; $\pi_{i}$ is comparing two policies, $\mathcal{D}^{i}_{s_{0},o_{0}}$ compares one policy under possible distribution shift. The utilized Type-I Error $\alpha^{*}$ and $N_{\text{max}}$ describe the constraints applied a priori by the evaluator (we underline to emphasize the change in $N_{\text{max}}$ for rows 1-3 and 4-6; observe that the sensitivity of the stopping times is very small). $N~{}\leq~{}N_{\text{max}}$ represents the amount of data available for the statistical analysis. We report the terminal empirical success rates (after $N$ trials) of each policy in each setting under $\hat{p}_{i}$ (this information is not available to any feasible algorithm). We do not have truth labels on this data; however, in all cases, every method arrived at the same decision, including the Oracle SPRT which has a priori access to privileged information $(\hat{p}_{0},\hat{p}_{1})_{N}$; this decision was Reject Null for all rows except the CarrotOnPlate tasks, which each returned Fail To Decide. We report the stopping times of all methods on the right of the table for every context; in all cases: lower is better. We put in bold any feasible method result that is near-optimal within ten trials (absolute) or 25% (relative) of the SPRT Oracle, which is not implementable by an evaluator. In the Multitask setting, we test $p_{1}>p_{0}$ uniformly across the preceding three tasks. This stopping time is the sum by column of the stopping times for the three tasks. Our method saves the evaluator over 200 trials in uniform certification over these three tasks as compared to either feasible baseline.
In CleanUpSpill, we compare the same policy on two different sets of initial conditions. The task is similar to the one originally presented by Xu et al. [60], which compares a set of ID initial conditions against the out-of-distribution (OOD) initial conditions. The ID set includes 10 initial conditions with a white towel and a short blue mug whereas the OOD set uses 10 with a checkered towel with a tall cyan mug (each initial condition is repeated five times). As shown in Table I (rows 4–6), the empirical gap of 52 percentage points was detected in 7–14 trials by all methods, though they were tuned (where applicable) to an $N_{\text{max}}$ up to thirty to seventy times larger; this demonstrates the significant reduction in sensitivity (from an evaluator’s standpoint) arising from setting $N_{\text{max}}$ versus choosing a batch size $N$. Furthermore, STEP’s efficiency only minimally degrades when $N_{\text{max}}$ is increased from 200 (row5) to 500 (row 6). In this setting, any of the three sequential methods would have prevented the need for nearly 70 of the 100 total rollouts (35 of the 50 batch trials per policy).
V-C2 Hardware Evaluation in the Small-Gap Regime
We evaluate STEP on two open-source vision-language-action (VLA) models: Octo-Base [40], an action-chunking transformer-based diffusion policy, and OpenVLA [27], an autoregressive policy leveraging a pretrained large language model backbone. All experiments were conducted in a toy kitchen environment from the Bridge Data V2 dataset [57], which is included in both policies’ training data. We considered the task of placing a carrot on a plate (denoted CarrotOnPlate, see Fig. 9 and Fig. 9), which is representative of evaluations investigated in [40, 27]. All policies were run on the Widow X 250S following the setup in [57].
For this task, the initial placement of the carrot is uniformly sampled from three possible locations (left, center, or right) on the counter, and the plate is placed next to the sink (see Figure 9). At the start of each trial, the robot joint angles are initialized such that the gripper is roughly aligned with the carrot initial position. The environment follows a categorical distribution with two outcomes: no object distractors or two object distractors. We consider two environment distributions: Env1 in which there are no distractors with probability $0.8$ and Env2 in which there are no distractors with probability $0.6$. Object distractors are sampled uniformly (without replacement) from the following object categories: orange, apple, green and blue sponges, brown and yellow cubes, eggplant, spoon, and towel. The initial locations of distractors is also sampled uniformly without replacement from four possibilities: on the stove, left of the stove, above the stove, or next to the faucet. The distractors are physically placed according to a uniform (continuous) distribution within the selected region. We sample 100 environment configurations each for three settings: i) Octo under distribution Env1, ii) OpenVLA under distribution Env1, and iii) Octo under distribution Env2. We then make the following comparisons with STEP: i) Octo (Env1) and Octo (Env2), and ii) Octo (Env1) and OpenVLA (Env1). In the first comparison (Table I, row 7), we test the effect of distribution shift in the probability of distractors being present. Here, the $\hat{p}_{0}$ corresponds to the perturbed distribution Octo (Env2), and $\hat{p}_{1}$ to the nominal distribution Octo (Env1). We find that, while there is an empirical gap (59% vs 68%) at $N=100$, no method returns a significant result at $\alpha^{*}=0.05$. In the second comparison (Table I, row 8), $\hat{p}_{0}$ corresponds to Octo (Env1) and $\hat{p}_{1}$ corresponds to OpenVLA (Env1). We observe no significant gap despite the empirical gap of 8 percentage points in favor of OpenVLA. Importantly, insignificance of the test does not mean that the null hypothesis should be accepted [20]; there remains a possibility that OpenVLA actually outperforms Octo, or that Octo’s performance indeed degrades due to the presence of distractors. However, our budget of $N_{\max}=100$ was likely not sufficient to accumulate enough evidence. In Section X-D in the Supplement, we further investigate this data insufficiency to show that, if the ground truth values were equal to the empirical success rates (59% vs. 68% and 68% vs. 76%), then we would require $N_{\max}=500$ trials to confidently determine $p_{1}>p_{0}$. This number is an order of magnitude larger than the current norms, reflecting fundamental yet often overlooked challenges in trustworthy policy comparison.
V-C3 Multi-Task Evaluation in SimplerEnv Simulation
Finally, we briefly consider the problem of multi-task and multi-policy extensions to this framework, and illustrate its potential via an example of policy evaluation in simulation (where costs are lower than on hardware, but still can be significant). Concretely, Octo-Small ($\pi_{1}$) and Octo-Base ($\pi_{0}$) [40] are compared in the SimplerEnv [32] simulation environment on three tasks of varying empirical difficulty. On the EggplantInBasket task, the policies each succeed at a near-50% rate, with a gap of 16.4 percentage points. For the SpoonOnTowel task, the gap is larger at 30 percentage points. For the hardest task, StackCube, the performance gap is 3 percentage points. The evaluation statistics are shown in Table I, rows 9-11. Note that the empirical success rates we observed are consistent with the findings of Li et al. [32] that Octo-Small is more performant on these tasks (see their Table V). We seek to evaluate the multitask comparison of the two policies. Letting $p_{s}^{[\tau]}$ denote the performance of Octo-Small and $p_{b}^{[\tau]}$ denote the performance of Octo-Base on task $\tau$, we test:
$\begin{split}&\mathbb{H}_{0}:\exists\tau\in\{1,2,3\}~{}p_{s}^{[\tau]}~{}\leq~{}p_{b}^{[\tau]}\\ &\mathbb{H}_{1}:\forall\tau\in\{1,2,3\}~{}p_{s}^{[\tau]}>p_{b}^{[\tau]}.\end{split}$ (11)
Many established and sophisticated methods exist to efficiently run multi-hypothesis testing (in this case, we are essentially evaluating three separate hypotheses, one for each task [^7] Note that multi-hypothesis testing can naturally handle the case of multi-policy comparison as well, where we would reduce the test to a set of pairwise comparisons that are examined simultaneously.). As a simple illustration, we use the standard Bonferroni (union bound) correction [15] to evaluate the test: specifically, running each of the individual three tests at level $\alpha^{*}_{[\tau]}=0.01$, we observe the stopping times shown in Table I, rows 9-11 (right-hand side). Via the Bonferroni correction, the combined decision (RejectNull, because every subtest decided RejectNull) expressed in LABEL:Eqn:MultiHypothesis is then confirmed at $\alpha=0.03=\sum_{\tau}\alpha^{*}_{[\tau]}$. As illustrated in Table I, each sequential method saves a substantial number of simulation rollouts on the easiest subtest (SpoonOnTowel). As expected, SAVI begins to struggle when the tests become more challenging, such as in EggplantInBasket. Finally, we observe the weakness of the Lai procedure: in heavily skewed cases, it suffers substantially even compared with SAVI methods, as shown in StackCube. To summarize: naive multitask evaluation requires the aggregation of multiple batches of rollouts, here totaling 1500 per policy (500 per task per policy). Note in rows 9-11 how different the number of requisite trials can be, and therefore how hard it is to reliably run the evaluation using a fixed batch size. On the easiest task, even when tuned to $N_{\text{max}}=500$, the comparison was answered in fewer than 40 rollouts by all sequential methods, a savings of over 90%. On the progressively harder cases the number of required samples increased 5-10 times over the easiest, but our method (STEP) improved substantially over each of the other sequential procedures. In total, STEP would have saved the evaluator an additional 200 rollouts for each of Octo-small and Octo-base for the multitask comparison problem as compared to the current SOTA approaches.
VI Related Work
Thus far, our discussion has focused solely on the practical setting of comparative policy evaluation. However, mathematical statistical testing has a well-established near-century-long history, and sequential testing in particular has existed for nearly the same amount of time. This section provides an extensive review of the statistics literature to further highlight the significance of our approach.
VI-A Statistical Testing and Policy Evaluation
The Neyman-Pearson (N-P) statistical testing paradigm [37] forms the foundation of the last century of frequentist statistical decision theory. Applicable to questions spanning many fields, these methods have generally been applied in robotics for predictable policy characterization [^8] As a simple example, one can accurately predict a priori that for estimating Ber($p$) with $\hat{p}\in[0.25,0.75]$ and $N~{}\geq~{}36$, a 95% confidence interval for $p$ will be approximately $\hat{p}\pm\frac{1}{\sqrt{N}}$. [53, 1], where the number of samples is fixed and a single decision or estimate is to be made. The Neyman-Pearson Lemma [37] and the Karlin-Rubin Theorem [25] give sufficient conditions for maximal power probability ratio tests (PRTs), which form the precursor to the SPRT Oracle used in this work. Specific methods have been developed for robot policy comparison-type problems in the context of 2x2 contingency tables. Of the tests by Fisher [17], Boschloo [6], and Barnard [3], the latter is a powerful batch procedure for comparison problems. However, while it has strong power in the batch setting, it does not provide a direct mechanism for choosing the appropriate size of the batch a priori.
VI-B Sequential Statistical Evaluation Methods
The difficulty in choosing the appropriate batch size motivates the sequential testing framework set out in Wald [55]. This is the setting adopted in this paper and described in Section III. Wald and Wolfowitz [56] showed that in the simple-vs-simple setting, the sequential probability ratio test (SPRT) minimizes the expected number of samples among all tests that control Type-I and Type-II error, extending the N-P result. Significant efforts have extended near-optimality results to the composite testing regime. Minimax results attempt to limit the worst-case expected sample size [26, 33, 16], while others minimize the expected sample size under a weighted mixture over the alternatives [47, 19]. Lai [30] reconciled this Bayesian interpretation with the frequentist developments of Chernoff [10, 11, 12].
VI-B1 Optimal Stopping-Based Methods
The direct approach to synthesizing near-optimal decision regions in the composite-vs-composite regime relies on developments in the theory of martingales and optional stopping [58]. Van Moerbeke [51] provides a clear reduction of the statistical decision making problem to that of optimal stopping and exposits previous developments in the area equating the optimal stopping procedure with the solution of a Stefan-type free-boundary partial differential equation (PDE) [8]. Unfortunately, while specification of the testing problem is usually feasible, its mapping to the PDE parameterization is generally difficult to specify under composite null hypotheses, rendering this method impractical. A parallel line of work considers utilizing asymptotic approximations of the solution to the free-boundary problem in order to construct near-optimal tests. The work of Lai [30] solves for a near-optimal procedure in the univariate composite-vs-composite setting and follow-on work [31, 9] extends this to the multivariate setting. While useful for proving asymptotic optimal rates, these methods suffer in the finite-$N_{\max}$ regime due to losing decision-making information.
VI-B2 Safe, Anytime-Valid Inference (SAVI) Methods
The difficulty in synthesizing optimal procedures in the multivariate setting (see, e.g., [31, 9]) motivates recent developments which fuse a distributionally robust safety invariance and the structure of the probability ratio test. This yields a family (SAVI) of methods which generalize to any $N_{\max}$ (i.e., arbitrarily small performance gaps) and a richer set of composite-hypothesis testing problems. Utilizing Ville’s Inequality (a sequential generalization of Markov’s Inequality) [52], SAVI methods construct a probability ratio test that enforces Type-I error control uniformly in time [22, 43]. In the small sample regime it is the effect of the Power-1 nature of the test [44, 29, 43] that can most negatively affect performance.
VI-C Numerical Implementations
Numerical methods are of significant practical importance in statistical testing. Excellent implementations of batch procedures have been released in the SciPy [54] package; recently, the particular problem of optimal binomial confidence intervals was addressed in [53]. However, despite a significant history dating back to the 1980s [13], numerical methods for sequential analysis in evaluation are relatively limited. Further, in [13] as in more recent work, emphasis has remained on the simple-null or univariate composite setting [38, 39, 16] for non-SAVI procedures. However, to our knowledge, a numerical implementation of near-optimal policy comparison procedures in the full multidimensional composite-vs-composite setting has not yet been implemented.
VII Limitations
Despite the strong performance of STEP as shown in Section V, it does have several practical and theoretical limitations. First, the development of STEP heavily relies on the mathematical structure of Bernoulli distributions, which essentially requires the policy evaluation results to be presented as binary success/failure metrics. Extending STEP to more complex discrete and continuous distributions would be a valuable future research direction for broader applicability. Second, in our current implementation a user needs to specify STEP’s risk accumulation rate $f(k)$. Although we empirically show that the uniform rate already achieves near-optimality, further research is needed to improve the performance.
Besides these limitations that are specific to STEP, researchers must be mindful of underlying assumptions and common misuse of statistical hypothesis testing [20]. Concretely, the foundational assumption is that evaluation data are i.i.d. as all the statistical assurances are built on top of it. To this end, we ensured that all hardware evaluations comported with the best practices of [28]. Finally, Goodhart’s Law [49] provides a useful warning about inadvertent p-hacking: having statistically rigorous evaluation is important, but if “significance at level $\alpha^{*}$” becomes a target and not a metric, it can induce undesirable research practices. We emphasize that p-values and significance levels are only as rigorous as the rigor of the research methods that utilize them. Their adoption has the potential to be enormously valuable for the empirical codification of robotics knowledge, but they are not a panacea.
VIII Future Work
There are several concrete extensions of STEP which promise significant benefits to practical robotic evaluation. Most direct is incorporation of more principled multi-policy evaluation, building upon Section V-C3. Additionally, partial credit schema have found increasing importance in long-horizon evaluation [28]; generalizing STEP beyond binary success/failure metrics to discrete (multinoulli) partial credit is direct in theory, although it presents additional challenges in implementation.
IX Conclusion
We present STEP, a novel sequential statistical method to rigorously compare performance of imitation learning policies through a series of evaluation trials. STEP’s sequential evaluation scheme provides flexibility in adapting the number of necessary trials to the underlying difficulty of policy comparison. This leads to sample efficiency in cases where one policy clearly outperforms the other, while avoiding overconfident and potentially incorrect evaluation decisions when the policies are closely competing. We show that STEP near-optimally minimizes the expected number of trials required in the policy comparison problem. Furthermore, STEP matches or exceeds the performance of state-of-the-art baselines across a wide swath of practical evaluation scenarios in numerical and robotic simulation and on numerous physical hardware demonstrations. These results highlight the practical utility of STEP as a versatile statistical analysis tool for policy comparison, contributing to the foundation of robot learning as an empirical science.
Acknowledgments
This work was primarily completed during an internship at Toyota Research Institute (TRI). The authors would like to thank Chen Xu for insightful feedback and discussions, as well as the robot teacher team at TRI headquarters for their data collection efforts, in particular Derick Seale and Donovan Jackson. Finally, we thank the anonymous reviewers and the area chair for incisive suggestions for improving the paper.